### v1

# Marangoni-driven pattern transition and the formation of ridges and hills in surfactant-covered parametric surface waves

Abstract
Nonlinear surface waves are excited via a parametric oscillation of a surfactant-covered interface. Increasing the relative magnitude of the surfactant-induced Marangoni stresses results in a pattern transition from squares (observed for surfactant-free interfaces) to asymmetric squares, weakly wavy stripes, and ridges and hills. These hills are a consequence of the bi-directional Marangoni stresses at the neck of the ridges. The mechanisms underlying the pattern transitions and the formation of exotic ridges and hills are discussed in this Letter.
Faraday [1] noticed that the vertical vibration of a fluid layer produces surface waves oscillating at half the driving frequency. Crossing a threshold amplitude, these Faraday surface waves usually organize into patterns like squares, hexagons, stripes, quasipatterns, and superlattices. Complications arise from factors such as contact line dissipation [2], multi-frequency forcing [3], and surface contamination [4]. Surfactant-covered Faraday waves have been studied to assess the damping rates of surface waves [5, 6, 7] due to the influence of Marangoni stresses [4, 8]. However, these studies are limited to linearised one- or two-dimensional models, with some being carried out using lubrication theory, lacking three-dimensional studies of strongly nonlinear Marangoni effects on pattern formation in Faraday waves.
In this Letter, we report the results of simulations of the three-dimensional surfactant-covered Faraday waves; we focus on the influence of Marangoni effects on the surface wave patterns. Our study reveals that increasing the relative importance of Marangoni stresses leads to transitions from square patterns to asymmetric squares, weakly wavy stripes, and ridges and hills. These ridges and hills are new features that occur at higher Marangoni number (or $B$ as defined in this Letter). Ridges are found to rise non-uniformly and fall by forming a hill. Our direct numerical simulations help uncover the rich physics of the dynamics of these newly observed ridges and hills.
Our computational domain is a laterally-periodic box of area $\tilde{\lambda}_{c}\times\tilde{\lambda}_{c}$ and height $\tilde{H}$, where $\tilde{\lambda}_{c}$ is the critical wavelength. A thin film of a water-glycerine mixture (of density $\tilde{\rho}_{w}=1000~{}\rm kg/m^{3}$, viscosity $\tilde{\mu}_{w}=0.025~{}\rm kg/m~{}s$) of depth $\tilde{h}=1~{}\rm{mm}$ is considered with air as the background fluid. The interface (of surface tension $\tilde{\sigma}=70~{}\rm mN/m$) is vibrated at a frequency $f=100~{}\rm{Hz}$ and acceleration amplitude $\tilde{A}$. We assume that a layer of insoluble surfactant monomers (of concentration $\tilde{\Gamma}$) covers the interface such that the number of moles of monomers on the surface is conserved. The surfactant-dependent surface tension is parameterised by the Langmuir-Szyszkowski equation of state, $\tilde{\sigma}(\tilde{\Gamma})/\tilde{\sigma}_{0}=1+\beta_{s}\ln(1-\tilde{\Gamma}/\tilde{\Gamma}_{\infty})$, where $\tilde{\Gamma}_{\infty}$ is the maximum packing concentration and $\beta_{s}$ is the elasticity parameter reflecting the effectiveness of the surfactant in lowering $\tilde{\sigma}$.
Figure 1: (a) Phase diagram in the $\beta_{s}-\Gamma_{0}$ parameter plane showing the inertia-dominated (violet) and Marangoni-dominated (pink) regions. The solid, dotted, dot-dashed, and dashed lines correspond to the $B=1,1.23,1.46,2$ contours, respectively. The four typical patterns are squares, asymmetric squares, weakly wavy stripes, and ridges and hills. The phase boundaries are accurate to within $\Delta B=\pm 0.1$. The corresponding values of $B$ and $F_{c}^{D}$ are reported in table II of the supplemental material [9]. (b) Spatiotemporal evolution of the surface deflection $\zeta$ over one time period is shown from left to right; squares ($B=0.92$), asymmetric squares ($B=1.23$), weakly wavy stripes ($B=1.51$), and ridges and hills ($B=1.83$) are shown from top to bottom rows, respectively. (c,d) $\zeta_{mn}$ and $\Gamma_{mn}$, the maximal magnitudes over time of the $\zeta$ and $\Gamma$ Fourier coefficients, respectively, as a function of $B$.
We initialize our simulations with a plane interface and uniformly distributed surfactant coverage $\tilde{\Gamma}_{0}$. The interfacial dynamics are coupled to a surfactant transport equation that accounts for inertia, viscosity, gravity, capillarity, Marangoni stresses, surface diffusion, and parametric forcing [10]. Non-dimensionalization of the equations is carried out, in which $\tilde{h}$, $1/2\pi f$, and $\tilde{\Gamma}_{\infty}$ are taken as the characteristic length, time, and concentration scales, leading to six dimensionless numbers. We set these and their values to be Reynolds $Re\equiv 2\pi\tilde{\rho}_{w}\tilde{h}^{2}f/\tilde{\mu}_{w}=25.3$, Weber $We\equiv 4\pi^{2}\tilde{\rho}_{w}\tilde{h}^{3}f^{2}/\tilde{\sigma}_{0}=18.37$, Péclet, $Pe\equiv 2\pi\tilde{h}^{2}f/\mathcal{D}=2.5\times 10^{5}$, and Froude $Fr\equiv 2\pi f\sqrt{\tilde{h}/g}=6.34$ numbers. Rather than the usual Marangoni number, $Ma=\beta_{s}/We$, we introduce $B\equiv(\tilde{\sigma}_{0}-\tilde{\sigma}(\tilde{\Gamma}_{0}))/(2\pi\tilde{\mu}_{w}\tilde{h}f)$, which is the ratio of vibratory inertial and Marangoni timescales, to quantify the combined effects of $\beta_{s}$ and $\tilde{\Gamma}_{0}$; $B$ is varied systematically to assess the emergent pattern transitions at the acceleration amplitude obtained by $Fr_{t}=2\pi f\sqrt{\tilde{A}/g}$. Detailed descriptions of the formulation, scaling, and numerical method [10] are presented in Section A of the supplemental material [9].
After a transient phase, Faraday waves appear, which correspond to subharmonic waves whose amplitude is steady and whose response period $T$ is twice that of the forcing period. We set $t=0$ to be an instant within the steady-amplitude Faraday-wave regime. Hereafter, we discuss our results in terms of dimensionless variables (with the tilde suppressed). The dimensionless critical wavelength $\lambda_{c}$ and acceleration amplitude $F_{c}=Fr/Fr_{t}$ for a clean interface are $5.30$ and $12.34$, respectively. We studied the effect of varying $B$ on the threshold acceleration $F_{c}^{D}$ with the details provided in Section B of the supplemental material [9]. Our results show that increasing $B$ stabilises the interface, as observed in previous studies [5, 4]. The computations for assessing the influence of $B$ on the interfacial dynamics in the nonlinear regime are then carried out for $F=1.1F_{c}^{D}$ for which square patterns are observed in the surfactant-free case.
As shown in Fig. 1(a), for $B<1$ (dark blue dots, purple region), the interface exhibits square symmetry. In a narrow band of $1\leq B\leq 1.23$ (light blue dots), the vertical and horizontal directions differ slightly; we refer to these patterns as asymmetric squares. Within $1.23\leq B\leq 1.46$ (orange dots), the asymmetric square pattern undergoes a transition to weakly wavy stripes. Ridges (ellipses whose major axes are in the $y$-direction) appear very faintly as dots for $B=1.23$, $t=3T/4$, and more prominently on the wavy stripes for $B=1.51$. For $B=1.83$, $t=0$, one can also see circular hills between each set of ridges. The hills are the continuation of the ridges formed in the previous half-period. One such instance is shown at $t=3T/4$, where the ridges have disappeared but the hills are present. We explore below the role of Marangoni stresses in the formation of these patterns.
Figure 2: (a)-(d) Three-dimensional visualization of the surface. (a) Rise of ridges and necking process at $t=T/4$ and (b) maximum rise of the ridge at $t=3T/8$. (c) Prominent hill on the ridge at $t=T/2$. (d) Falling hill at $t=3T/4$. (e-h) Two-dimensional projections on $x-z$ slice containing interface curve $s$ (indicated in (a)) for $t=T/4,3T/8,T/2,$ and $43T/$ 4, respectively. A half-wavelength (ridge to trough) is shown. Color-coding of the plane indicates $y-$vorticity $\omega_{y}$, while streamlines show flow in $x-z$ plane. The interface curve $s$ is colored according to the surfactant concentration. Red dots indicate the point of maximum curvature. (i-l) Tangential (see (f)) Marangoni stress and velocity along $s$ at $t=T/4,3T/8,T/2,$ and $43T/$ 4 shown as black and red curves, respectively. When the sign of one of these quantities is positive (negative), its direction points rightwards (leftwards) from the apex (trough) through the neck to the trough (apex) of the ridge, as indicated at the top (bottom) of Fig. 2(i). The vertical dashed line indicates the necking region, shown as the red dot in the corresponding $x-z$ projection. The length of $s$ decreases from about $5$ at $t=T/4,3T/8,$ $T/2$ to about $4$ at $t=3T/4$, as can be seen in the curves in (e-h).
To quantify the patterns, we evaluate the spatial Fourier spectra for the surface height, $\zeta$, and surfactant concentration, $\Gamma$, defining $\hat{\zeta}_{mn}(t)$ and $\hat{\Gamma}_{mn}(t)$ to be the Fourier coefficients associated with the $(x,y)$ wavevector $\mathbf{k}_{mn}$. We then set $\zeta_{mn}\equiv\max_{\left[t,t+T\right]}{|\hat{\zeta}_{mn}(t)|}$ and $\Gamma_{mn}\equiv\max_{\left[t,t+T\right]}{|\hat{\Gamma}_{mn}(t)|}$. Figures 1(c,d) present an overview of the spatial Fourier spectra of $\zeta$ and $\Gamma$ as a function of $B$ in the range $B\in[0,1.51]$. At higher $B$, ridges and hills emerge, where steep spatial gradients and many higher spatial harmonics appear.
For $B<1$, the square pattern is characterized by comparable amplitudes of $\zeta_{10}$ and $\zeta_{01}$, as shown in Fig. 1(c). For $B<0.5$, where Marangoni effects are weak, the $\zeta_{mn}$ modes have magnitudes similar to those associated with the clean case corresponding to $B=0$, consistent with previous findings [11, 12]. For $B>1$, Marangoni-driven stresses dominate over inertial effects. The square symmetry is broken, and by $B=1.23$, $\zeta_{10}$ surpasses $\zeta_{01}$, with an increase in higher-order modes, such as the $\zeta_{20}$ mode. As $B$ increases further, strong $x-$dependent modes emerge, leading to a transition from asymmetric squares to stripes (see Fig. 1(c)).
A parallel change occurs in the $\Gamma$-spectrum. For $B<1$, the surfactant is advected without being significantly hindered by Marangoni stresses, aligning the $\Gamma$-spectrum with the $\zeta$-spectrum, where $\Gamma_{10}$ and $\Gamma_{01}$ dominate (see Fig. 1(d)). For $B>1$, $\Gamma_{10}$ begins to surpass $\Gamma_{01}$. Thus, $B\approx 1$ is a pivotal point in the dynamics, at which there is an equilibrium between the opposing mechanisms of advection-driven surfactant inhomogeneity and Marangoni-driven homogeneity.
We now turn to the formation of hills and ridges on the interface. Figures 2(a-d) illustrate the evolution of a small portion of the interface, color-coded by surfactant concentration. During the first half-cycle, the ridges rise, and the fluid and surfactant flow up from the troughs, advecting the surfactant to the apex of the ridge. Figures 2(e-h) show two-dimensional projections containing arc $s$, as indicated in Fig. 2(a). As the surfactant is advected towards the apex, a $\Gamma$-deficit (higher $\sigma$) is created at the trough.
The capillary force resulting from the $\Gamma$-deficit leads to the emergence of a bulb on the ridge, surrounded by a narrow region of negative curvature, which we call a neck, and which is highlighted by a red spot on the interface in Figs. 2(e-h); $\Gamma$ accumulates at the ends of the ridge as shown in Figs. 2(a,b). Marangoni stresses along $s$ counteract the $\Gamma$-inhomogeneity caused by the surface advection. This is shown in Fig. 2(i), where $\tau>0$ and $u_{t}<0$ along the arc $s$. We call this a barrier. This barrier rigidifies the surface during the first half-cycle, leading to $|\mathbf{u}_{s}|\approx 0$ at $t=3T/8$, as shown in Fig. 2(b).
The negative vorticity along the surface in Fig. 2(f) indicates that $\tau$ opposes the surface advection. Due to this barrier, a backflow develops on the surface from the apex towards the neck, as indicated by $u_{t}>0$ in the inset of Fig. 2(j). This drives surfactants from the apex towards the neck. Simultaneously, the accumulated surfactant at the ends of the ridge flows towards the neck due to a similar mechanism, as illustrated by the red arrows in Fig. 2(b). During this process, the midpoint of the ridge rises to form a bulb; see Fig. 2(c). By $t=T/2$, $\Gamma$ is maximal (so $\sigma$ is minimal) at the neck.
The accumulated surfactant causes Marangoni stresses, with distinct peaks of $\tau>0$ and $\tau<0$ across the neck (Fig. 2(k)). The barrier is now formed at the neck (shown as a white dotted region in Fig. 2(c)) where these stresses in the region between the apex and the neck begin to oppose the flow reversal at half-cycle. Meanwhile, surface tension decreases at the neck. As a result, the neck begins to reopen (see the streamlines in fig. 2(g)) as is commonly observed in surfactant-laden neck reopening phenomena, discussed in detail in [13, 12].
[…]
Figure 3: Marangoni-influenced ridge formation: (a) $x-z$ projection containing $s$, as defined in Fig. 2, at $t=T$; the color-coding used here is that of Fig. 2. (b) Three-dimensional visualization of the interface color-coded by the magnitude of Marangoni stresses $|\nabla_{s}\sigma|$, indicating the barriers around the rising ridge.
Figure 3 further elucidates the mechanism of ridge formation. The surfactant accumulates on the developing ridge due to a combination of Marangoni-driven surface flow from the neck to the trough ($\tau>0$ in Fig. 2(h,l)) as previously discussed, and advection through bulk flow in the second half-cycle leading to strong surface compression at the ridge. This accumulation (see $\Gamma$-surplus region highligted in Fig. 3(a)) generates a Marangoni stress, directed from the newly developed ridge toward the falling hill (as highlighted by the arrow indicating the direction of $\tau$ in Fig. 3(a)). The magnitude of the Marangoni stress, $|\nabla_{s}\sigma|$, is shown in Fig. 3(b). This high-stress region, which surrounds the developing ridge, highlights the strength of the barrier to ridge formation. Close inspection of this region reveals that the barrier is weaker at the midpoint of the ridge, allowing stronger inward-directed surface flow to this region (viz. the velocity glyphs in Fig. 3(b)). This, in turn, leads to a higher elevation at the midpoint of the ridge than at its ends, as shown in Fig. 3(b).
The study highlights the role of Marangoni stresses in Faraday wave pattern transitions. Using a dimensionless parameter $B$ to compare the Marangoni and inertial timescales, we found that for $B\approx 1$, the square patterns are succeeded by asymmetric squares and then by weakly wavy stripes as $B$ is further increased. The novel finding highlighted here is that for higher $B$ values, ridges and hills appear. Due to strong Marangoni flow during a cycle of forcing, surfactant- and vibration-induced surface flows compete, giving rise to a Marangoni-driven barrier, which is the cause of these interesting ridges and hills.
We thank EPSRC for support through the PREMIERE (EP/T000414/1) programme grant, and the ANTENNA Prosperity Partnership (EP/V056891/1). D.J. and J.C. thank the Institut du Developpement et des Ressources en Informatique Scientifique of the CNRS, coordinated by GENCI (Grand Equipement National de Calcul Intensif) grant 2024 A0162B06721. D.P. thanks Imperial for a President PhD scholarship, and Prof. Cédric Beaume for fruitful discussions.